If fixed costs are $251,000, the unit selling price is $33, and the unit variable costs are $18, the break-even sales (units) if fixed costs are reduced by $30,500 is

a.17,640 units

b.22,050 units

c.14,700 units

d.11,760 unit
Answer #1

Contribution margin=Sales-Variable cost

=33-18=$15 per unit

New fixed costs=251,000-30500=$220500

Hence breakeven=Fixed cost/contribution margin

=220500/15

=14700 units.